# Are queries and keys always relevant? A case study on Transformer wave functions

### v1

@@ Abstract
The dot product attention mechanism, originally designed for natural language processing (NLP) tasks, is a cornerstone of modern Transformers. It adeptly captures semantic relationships between word pairs in sentences by computing a similarity overlap between queries and keys. In this work, we explore the suitability of Transformers, focusing on their attention mechanisms, in the specific domain of the parametrization of variational wave functions to approximate ground states of quantum many-body spin Hamiltonians. Specifically, we perform numerical simulations on the two-dimensional $J_{1}$-$J_{2}$ Heisenberg model, a common benchmark in the field of quantum-many body systems on lattice. By comparing the performance of standard attention mechanisms with a simplified version that excludes queries and keys, relying solely on positions, we achieve competitive results while reducing computational cost and parameter usage. Furthermore, through the analysis of the attention maps generated by standard attention mechanisms, we show that the attention weights become effectively input-independent at the end of the optimization. We support the numerical results with analytical calculations, providing physical insights of why queries and keys should be, in principle, omitted from the attention mechanism when studying large systems. Interestingly, the same arguments can be extended to the NLP domain, in the limit of long input sentences.
@@ 1 Introduction
footnotetext: *Equal contribution. Author ordering is random.
Transformers [1] have emerged as one of the most powerful deep learning tools in recent years. They are task-agnostic neural networks that are pre-trained to build context-sensitive representations of words in input sentences [2, 3, 4]. The success of Transformers lies in their remarkable flexibility: with minimal modifications, they excel in addressing diverse problem domains, often outperforming specialized approaches [5, 6, 7]. This is a consequence of their versatile foundational components, namely dot product Self-attention, Multilayer perceptron (MLP), Layer Normalization, and skip connections. While elements like the MLP, Layer Normalization, and skip connections are task-agnostic and offer broad applicability, the functional form of the dot product attention mechanism was originally tailored for NLP tasks. In this context, a sentence is processed by initially associating each word with a vector through a lookup table. These vectors form a sequence which is processed by the Self-attention Mechanism [1], designed to generate, for each input, an output vector as a weighted sum of all other inputs. Crucially, the coefficients in this sum involve learnable parameters that are optimized to capture the semantic relationships between pairs of words within the sentence. The remarkable generalization properties of Transformers in NLP tasks has been associated to the use of attention weights that depend on the input values, thereby capturing powerful inductive biases related to the semantics in natural languages [8, 9]. One wonders if the dot product attention mechanism provides an inductive bias which is the most appropriate in any data domain. For example, in the context of protein contact prediction, Ref. [10] suggests that input-independent attention weights achieve competitive performance compared to the standard approach. In this paper, we delve into this aspect by exploring the application of the Transformer architecture as a Neural-Network Quantum State (NQS) for approximating the ground state of quantum many-body spin Hamiltonians on lattice [11]. The Transformer architecture has already been employed in this context, achieving highly accurate results across different systems [12, 13, 14, 15, 16, 17, 18, 19]. While many of these works adopt the standard attention mechanism [13, 16, 17], Ref. [14] employ a simplified version, omitting queries and keys, still reaching state-of-the-art accuracy on one of the most popular benchmark problem in frustrated magnetism. Therefore, the question whether queries and keys provide a suitable inductive bias for general applications persists. In this work, we tackle this question by systematically investigating the performance of different attention mechanism within Transformer wave functions. In the following we summarized our main findings:
@@
    In Transformer wave functions, the standard dot product attention mechanism used in NLP does not improve the performance of a simpler mechanism in which the attention weights are input-independent.
@@ 2.1 The quantum-many body problem
The physical properties of an interacting quantum-many body system described by an Hamiltonian $\hat{H}$ are determined by solving the time-independent Scrödinger equation $\hat{H}\ket{\Psi_{n}}=E_{n}\ket{\Psi_{n}}$, where |$\ket{\Psi_{n}}$ and $E_{n}$ are eigenstates and eigenvalues of $\hat{H}$, respectively. In principle, fixing a basis in the Hilbert space, we can numerically obtain the spectrum of $\hat{H}$ by storing all its matrix elements and using standard computational routines to diagonalize it. However, a critical challenge arises due to the exponential growth in the size of this matrix with respect to the number of particles in the system, rendering this approach feasible only for small systems [20]. Typically, the focus lies in the low-energy properties of the Hamiltonian, particularly in its ground state |$\ket{\Psi_{0}}$. To obtain approximations of the ground state for systems where exact diagonalization is not feasible, many methods have been developed over the years. Here, we focus on variational approaches, where a variational state |$\ket{\Psi_{\theta}},$ depending on a set of $P$ parameters $\theta$, is optimized to minimize the variational energy $E_{\theta}=\braket{\Psi_{\theta}}{\hat{H}}{\Psi_{\theta}}/\braket{\Psi_{\theta}}{\Psi_{\theta}}$. According to the Variational Principle [21], the energy $E_{\theta}$ associated to any generic state |$\ket{\Psi_{\theta}}$ is always bigger then the ground state energy $E_{\theta}\leq E_{0}$. Moreover, provided that the ground state is unique, we have that $E_{\theta}=E_{0}$ if and only if |$\ket{\Psi_{\theta}}=\ket{\Psi_{0}}$. To be concrete, we consider systems of $N$ spin-$1/2$ arranged on regular lattices. In this case, the variational state can be expanded as |$\ket{\Psi_{\theta}}=\sum_{\{\sigma\}}\Psi_{\theta}(\sigma)\ket{\sigma}$, where {|$\{\ket{\sigma}=\ket{\sigma_{1}^{z},\sigma_{2}^{z},\dots,\sigma_{N}^{z}}\}$ with $\sigma^{z}_{i}=\pm 1$ is the computational basis. The many-body wave function $\Psi_{\theta}(\sigma)=\braket{\sigma}{\Psi_{\theta}}$ is a compact representation of the quantum state, which maps each configuration of the basis set |$\ket{\sigma}$ to a complex number using a relatively small number of parameters $P$ compared to the exponential size of the full Hilbert space ($P\ll 2^{N}$).
@@ 2.2 Variational Monte Carlo Framework
5:     Stochastic estimation of the gradient of the energy : $F_{\alpha}=-\partial_{\alpha}E_{\theta}$ with $\alpha=1,\dots,P$
6:     Stochastic estimation of the Quantum Geometric Tensor : $S_{\alpha,\beta}$ with $\alpha,\beta=1,\dots,P$
7:     Update of the parameters with Stochastic Reconfiguration: $\delta\theta_{\alpha}=\tau\sum_{\beta}S^{-1}_{\alpha,\beta}F_{\beta}$
@@
$F_{\alpha}=-\frac{\partial E_{\theta}}{\partial\theta_{\alpha}}=-2\Re\left[\braket{(\hat{H}-\braket{\hat{H}})(\hat{O}_{\alpha}-\braket{\hat{O}_{\alpha}})}\right]\ ,$ (1)
where $\alpha=1,\dots,P$ and $\hat{O}_{\alpha}$ is a diagonal operator defined as ${O_{\alpha}(\sigma)={\partial}\text{Log}[\Psi_{{\theta}}(\sigma)]/{\partial\theta_{\alpha}}}$. The latter log-derivative can be efficiently computed for NQS architectures using automatic differentiation [23]. The expectation values $\braket{..}$ with respect to the variational state |$\ket{\Psi_{\theta}}$ are stochastically estimated using Markov Chain Monte Carlo (MCMC), where spin configurations are sampled according to $|\Psi_{\theta}(\sigma)|^{2}$ [22]. The parameters are updated according to the Stochastic Reconfiguration (SR) method [24, 25] (see step 7 in Algorithm 1), which is formally equivalent to Natural Gradient [26, 27]. The SR approach takes into account the geometric properties of the energy landscape through the Quantum Geometric Tensor $S$, a $P\times P$ matrix which generalizes the Fisher information metric [28]:
$S_{\alpha,\beta}=\Re\left[\braket{(\hat{O}_{\alpha}-\braket{\hat{O}_{\alpha}})^{\dagger}(\hat{O}_{\beta}-\braket{\hat{O}_{\beta}})}\right]\ .$ (2)
Recent studies have demonstrated the effectiveness of SR in optimizing NQS with a large number of parameters [29, 14, 15]. It is important to stress that in VMC the data, i.e., spin configurations, are generated on the fly during the optimization process by sampling from $|\Psi_{\theta}(\sigma)|^{2}$. This is different from conventional machine learning scenarios where a fixed training set is provided.
@@ 2.3 Vision Transformer wave function
In 2017, Carleo and Troyer [11] proposed to use neural networks to parametrize the variational quantum state amplitudes $\Psi_{\theta}(\sigma)\in\mathbb{C}$. Neural-Network Quantum States (NQS) have demonstrated remarkable representational power in challenging problems [30, 31] and reached the state-of-the-art results in describing the ground state properties of two-dimensional frustrated magnets [32, 14, 15, 29, 33], bosonic [34] and fermionic [35, 36, 37, 38] models. In this work, we focus on a particular NQS based on the Vision Transformer (ViT) architecture, introduced in Ref. [15]. First, following what is done for images [7], each input spin configuration $\sigma$ is splitted into $n$ patches, which are linearly embedded in a $d$-dimensional space, thus producing a sequence of vectors $(\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n})$, with $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$. This sequence is processed by a deep ViT with real-valued parameters which produces an output sequence of vectors $(\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{n})$, with $\boldsymbol{y}_{i}\in\mathbb{R}^{d}$. The ViT architecture is constituted by $n_{l}$ encoder blocks, each of them including Multi-Head attention with $h$ heads, two-layers MLP with GeLU activation, skip connections and Pre-Layer Normalization [39]. Then, a $d$-dimensional hidden representation is obtained as $\boldsymbol{z}=\sum_{i=1}^{n}\boldsymbol{y}_{i}\in\mathbb{R}^{d}$. Only at the end, the latter is mapped to a complex number representing the logarithm of the amplitude. This final mapping is performed by an output layer parametrized as a shallow network, namely $\text{Log}[\Psi_{\theta}(\sigma)]=\sum_{\alpha=1}^{d}g(b_{\alpha}+\boldsymbol{w}_{\alpha}\cdot\boldsymbol{z})$, with non-linearity $g(\cdot)=\text{logcosh}(\cdot)$ and complex-valued trainable parameters $\{b_{\alpha},\boldsymbol{w}_{\alpha}\}_{\alpha=1}^{d}$. For more details about the architecture see Ref. [15].
@@ 3.1 Relative positional attention mechanisms
The attention weights $\alpha_{ij}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ form a $n\times n$ matrix and measure the relative importance of the $j$-$th$ input when computing the new representation of the $i$-$th$ input.
@@
    Factored attention, introduced in Refs. [10, 42]:
@@
The vectors $Q\boldsymbol{x}_{i}$, $K\boldsymbol{x}_{i}$ and $V\boldsymbol{x}_{i}$ are called queries, keys and values, respectively. The matrices $Q$, $K$ and $V$, along with the positional encoding $P$, are trainable parameters. In Fig. 1, we show a schematic representation of these three different attention mechanisms. The Factored version has a reduced number of parameters, being the attention weights input independent. Regarding the computational cost for the calculation of each attention weight, we have $O(1)$ complexity in the Factored case and $O(nd^{2})+O(n^{2}d)$ in the other two cases. Decoupled attention, as represented by Eq. (5), is the simplest extension of the Factored version in Eq. (6), where the attention weights now factor in the input dependence: setting $Q=K=0$ allows to recover the Factored attention, albeit with a constant shift. Instead, in T5 attention [see Eq. (4)] all the attention weights are constrained to be positive due to the global softmax activation.
@@ 4 Results
Figure 2: Relative error $\Delta\varepsilon=|(E_{0}-E_{\text{ViT}})/E_{0}|$ during the optimization of the ViT wave function on the $J_{1}$-$J_{2}$ Heisenberg model at $J_{2}/J_{1}=0$ (left panel) and at $J_{2}/J_{1}=0.5$ (right) on a ${6\times 6}$ lattice with periodic boundary conditions. The exact energies $E_{0}$ are computed with exact-diagonalization approaches. All architectures are trained with the same optimization protocol, using SR (see section 2.2) with cosine decay scheduler for the learning rate with an initial value of $\tau=0.03$. While increasing the number of parameters in the neural network enhances accuracy, as evidenced by considering four (panel b-d) instead of one (panel a-c) layer architectures, our simulations reveal that both T5 and Decoupled attention mechanisms, despite the increased number of parameters and computational complexity, yield comparable results to those of Factored attention. The optimization curves are consistent across multiple runs with different random initialization of the parameters.

### v2

@@ Abstract
The dot product attention mechanism, originally designed for natural language processing tasks, is a cornerstone of modern Transformers. It adeptly captures semantic relationships between word pairs in sentences by computing a similarity overlap between queries and keys. In this work, we explore the suitability of Transformers, focusing on their attention mechanisms, in the specific domain of the parametrization of variational wave functions to approximate ground states of quantum many-body spin Hamiltonians. Specifically, we perform numerical simulations on the two-dimensional $J_{1}$-$J_{2}$ Heisenberg model, a common benchmark in the field of quantum many-body systems on lattice. By comparing the performance of standard attention mechanisms with a simplified version that excludes queries and keys, relying solely on positions, we achieve competitive results while reducing computational cost and parameter usage. Furthermore, through the analysis of the attention maps generated by standard attention mechanisms, we show that the attention weights become effectively input-independent at the end of the optimization. We support the numerical results with analytical calculations, providing physical insights of why queries and keys should be, in principle, omitted from the attention mechanism when studying large systems.
@@ 1 Introduction
footnotetext: *Equal contribution.
Transformers [1] have emerged as one of the most powerful deep learning tools in recent years. They are task-agnostic neural networks that are pre-trained to build context-sensitive representations of words in input sentences [2, 3, 4]. The success of Transformers lies in their remarkable flexibility: with minimal modifications, they excel in addressing diverse problem domains, often outperforming specialized approaches [5, 6, 7]. This is a consequence of their versatile foundational components, namely the dot product self-attention mechanism, Multilayer Perceptron (MLP), Layer Normalization, and skip connections. While elements like the MLP, Layer Normalization, and skip connections are task-agnostic and offer broad applicability, the functional form of the dot product attention mechanism was originally tailored for natural language processing (NLP) tasks. In this context, a sentence is processed by initially associating each word with a vector through a lookup table. These vectors form a sequence which is processed by the self-attention mechanism [1], designed to generate, for each input, an output vector as a weighted sum of all other inputs. Crucially, the coefficients in this sum involve learnable parameters that are optimized to capture the semantic relationships between pairs of words within the sentence. The remarkable generalization properties of Transformers in NLP tasks have been associated with the use of attention weights that depend on the input values, thereby capturing powerful inductive biases related to the semantics in natural languages [8, 9]. One wonders if the dot product attention mechanism provides an inductive bias which is the most appropriate in any data domain. For example, Ref. [10] in the context of protein contact prediction and Ref. [11] in computer vision tasks suggest that input-independent attention weights achieve competitive performance compared to the standard approach. In this paper, we delve into this aspect by exploring the application of the Transformer architecture as a Neural-Network Quantum State (NQS) for approximating the ground state of quantum many-body spin Hamiltonians on lattice [12]. The Transformer architecture has already been employed in this context, achieving highly accurate results across different systems [13, 14, 15, 16, 17, 18, 19, 20]. While many of these works adopt the standard attention mechanism [14, 17, 18], Ref. [15] employs a simplified version, omitting queries and keys, still reaching state-of-the-art accuracy on one of the most popular benchmark problems in frustrated magnetism. Therefore, the question of whether queries and keys provide a suitable inductive bias for general applications persists. In this work, we tackle this question by systematically investigating the performance of different attention mechanisms within Transformer wave functions. In the following, we summarize our main findings:
@@
    In Transformer wave functions, the standard dot-product attention mechanism used in NLP does not improve the performance of a simpler mechanism in which the attention weights are input-independent.
@@ 2.1 The quantum-many body problem
The physical properties of an interacting quantum-many body system described by a Hamiltonian $\hat{H}$ are determined by solving the time-independent Scrödinger equation $\hat{H}\ket{\Psi_{n}}=E_{n}\ket{\Psi_{n}}$, where |$\ket{\Psi_{n}}$ and $E_{n}$ are eigenstates and eigenvalues of $\hat{H}$, respectively. In principle, fixing a basis in the Hilbert space, we can numerically obtain the spectrum of $\hat{H}$ by storing all its matrix elements and using standard computational routines to diagonalize it. However, a critical challenge arises due to the exponential growth in the size of this matrix with respect to the number of particles in the system, rendering this approach feasible only for small systems [21]. Typically, the focus lies in the low-energy properties of the Hamiltonian, particularly in its ground state |$\ket{\Psi_{0}}$. To obtain approximations of the ground state for systems where exact diagonalization is not feasible, many methods have been developed over the years. Here, we focus on variational approaches, where a variational state |$\ket{\Psi_{\theta}},$ depending on a set of $N_{p}$ parameters $\theta$, is optimized to minimize the variational energy $E_{\theta}=\braket{\Psi_{\theta}}{\hat{H}}{\Psi_{\theta}}/\braket{\Psi_{\theta}}{\Psi_{\theta}}$. According to the Variational Principle [22], the energy $E_{\theta}$ associated to any generic state |$\ket{\Psi_{\theta}}$ is always bigger than the ground state energy $E_{\theta}\geq E_{0}$. Moreover, provided that the ground state is unique, we have that $E_{\theta}=E_{0}$ if and only if |$\ket{\Psi_{\theta}}=\ket{\Psi_{0}}$. To be concrete, we consider systems of $N$ spin-$1/2$ arranged on regular lattices. In this case, the variational state can be expanded as |$\ket{\Psi_{\theta}}=\sum_{\{\sigma\}}\Psi_{\theta}(\sigma)\ket{\sigma}$, where {|$\{\ket{\sigma}=\ket{\sigma_{1}^{z},\sigma_{2}^{z},\dots,\sigma_{N}^{z}}\}$ with $\sigma^{z}_{i}=\pm 1$ is the computational basis. The many-body wave function $\Psi_{\theta}(\sigma)=\braket{\sigma}{\Psi_{\theta}}$ is a compact representation of the quantum state, which maps configurations of the basis set |$\ket{\sigma}$ to complex numbers using a relatively small number of parameters $N_{p}$ compared to the exponential size of the full Hilbert space ($N_{p}\ll 2^{N}$).
@@ 2.2 Variational Monte Carlo Framework
5:     Stochastic estimation of the gradient of the energy : $F_{\gamma}=-\partial_{\gamma}E_{\theta}$ with $\gamma=1,\dots,N_{p}$
6:     Stochastic estimation of the Quantum Geometric Tensor : $S_{\gamma,\beta}$ with $\gamma,\beta=1,\dots,N_{p}$
7:     Update of the parameters with Stochastic Reconfiguration: $\delta\theta_{\gamma}=\tau\sum_{\beta}S^{-1}_{\gamma,\beta}F_{\beta}$
@@
$F_{\gamma}=-\frac{\partial E_{\theta}}{\partial\theta_{\gamma}}=-2\Re\left[\braket{(\hat{H}-\braket{\hat{H}})(\hat{O}_{\gamma}-\braket{\hat{O}_{\gamma}})}\right]\ ,$ (1)
where $\gamma=1,\dots,N_{p}$ and $\hat{O}_{\gamma}$ are diagonal operators defined as ${O_{\gamma}(\sigma)={\partial}\text{Log}[\Psi_{{\theta}}(\sigma)]/{\partial\theta_{\gamma}}}$. The latter log-derivative can be efficiently computed for NQS architectures using automatic differentiation [24]. The expectation values $\braket{\dots}$ are are stochastically estimated using Markov Chain Monte Carlo (see Appendix A) by sampling $M$ configurations according to the amplitudes $|\Psi_{\theta}(\sigma)|^{2}$ (details can be found in Appendix B). The parameters are updated according to the Stochastic Reconfiguration (SR) method [25, 26] (see step 7 in Algorithm 1), which is formally equivalent to Natural Gradient [27, 28]. The SR approach takes into account the geometric properties of the energy landscape through the Quantum Geometric Tensor $S$, a $P\times P$ matrix which generalizes the Fisher information metric [29]:
$S_{\gamma,\beta}=\Re\left[\braket{(\hat{O}_{\gamma}-\braket{\hat{O}_{\gamma}})^{\dagger}(\hat{O}_{\beta}-\braket{\hat{O}_{\beta}})}\right]\ .$ (2)
Recent studies have demonstrated the effectiveness of SR in optimizing NQS with a large number of parameters [30, 15, 16]. It is important to stress that in VMC the data, i.e., spin configurations, are generated “on the fly” during the optimization process by sampling from $|\Psi_{\theta}(\sigma)|^{2}$. This is different from conventional machine learning scenarios where a fixed training set is provided.
@@ 2.3 Vision Transformer wave function
In 2017, Carleo and Troyer [12] proposed using neural networks to parametrize the variational quantum state amplitudes $\Psi_{\theta}(\sigma)\in\mathbb{C}$. Neural-Network Quantum States have demonstrated remarkable representational power in challenging problems [31, 32] and reached state-of-the-art results in describing the ground state properties of two-dimensional frustrated magnets [33, 15, 16, 30, 34], bosonic [35] and fermionic [36, 37, 38, 39] models. In this work, we focus on a particular NQS based on the Vision Transformer (ViT) architecture, introduced in Ref. [16]. First, following what is done for images [7], each $L\times L$ input spin configuration $\sigma$ is split into patches of size $b\times b$, which are linearly embedded in a $d$-dimensional space, thus producing a sequence of $n=L^{2}/b^{2}$ vectors $(\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n})$, with $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$. This sequence is processed by a deep ViT with real-valued parameters that produce an output sequence of vectors $(\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{n})$, with $\boldsymbol{y}_{i}\in\mathbb{R}^{d}$. The ViT architecture is constituted by $n_{l}$ encoder blocks, each of them including Multi-Head attention with $h$ heads, two-layer MLP with GeLU activation, skip connections and Pre-Layer Normalization [40]. Then, a $d$-dimensional hidden representation is obtained as $\boldsymbol{z}=\sum_{i=1}^{n}\boldsymbol{y}_{i}\in\mathbb{R}^{d}$. Only at the end, the latter is mapped to a complex number representing the logarithm of the amplitude. This final mapping is performed by an output layer parametrized as a shallow network, namely $\text{Log}[\Psi_{\theta}(\sigma)]=\sum_{\beta=1}^{d}g(b_{\beta}+\boldsymbol{w}_{\beta}\cdot\boldsymbol{z})$, with non-linearity $g(\cdot)=\text{logcosh}(\cdot)$ and complex-valued trainable parameters $\{b_{\beta},\boldsymbol{w}_{\beta}\}_{\beta=1}^{d}$. For more details about the architecture see Ref. [16].
@@ 3.1 Relative positional attention mechanisms
The attention weights $\alpha_{ij}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ form a $n\times n$ matrix, where $n$ is the number of patches, which measure the relative importance of the $j$-$th$ input when computing the new representation of the $i$-$th$ input.
@@
    Factored attention, introduced in Refs. [11, 10, 43]:
@@
The vectors $Q\boldsymbol{x}_{i}$, $K\boldsymbol{x}_{i}$ and $V\boldsymbol{x}_{i}$ are called queries, keys and values, respectively. The matrices $Q$, $K$ and $V$, along with the positional encoding $P$, are trainable parameters. When using relative positional encoding, the matrix $P$ is a circulant matrix with dimensions $n\times n$ which is constructed by different circular shifts of a vector of parameters in different rows. This results in only $n$ independent trainable parameters, denoted by $p_{i-j}$. In Fig. 1, we show a schematic representation of these three different attention mechanisms. The Factored version has a reduced number of parameters, being the attention weights input independent. Regarding the computational cost for the calculation of each attention weight, we have $O(1)$ complexity in the Factored case and $O(nd^{2})+O(n^{2}d)$ in the other two cases. Decoupled attention, as represented by Eq. (5), is the simplest extension of the Factored version in Eq. (6), where the attention weights now factor in the input dependence: setting $Q=K=0$ allows recovering the Factored attention, albeit with a constant shift. Instead, in T5 attention [see Eq. (4)] all the attention weights are constrained to be positive due to the global softmax activation.
@@ 4 Results
Figure 2: Relative error $\Delta\varepsilon=|(E_{0}-E_{\text{ViT}})/E_{0}|$ during the optimization of the ViT wave function on the $J_{1}$-$J_{2}$ Heisenberg model at $J_{2}/J_{1}=0$ (left panel) and at $J_{2}/J_{1}=0.5$ (right) on a ${6\times 6}$ lattice with periodic boundary conditions. The exact energies $E_{0}$ are computed with exact-diagonalization approaches. The architectures used for the simulations have $h=10$ heads, embedding dimension $d=60$, linear patch size $b=2$, $n_{l}=1$ layer in panels (a),(c), and $n_{l}=4$ layers in panels (b),(d). All networks are trained with the same optimization protocol, using SR (see section 2.2) for $5\times 10^{3}$ optimization steps with $M=6\times 10^{3}$ samples for the stochastic estimates. Each optimization step corresponds to one iteration in the for loop in Algorithm 1. A cosine decay learning rate scheduler is applied, starting with an initial value of $\tau=0.03$. The optimization curves are consistent across multiple runs with different random initialization of the parameters.